With the probability of .10, if we randomly selected 3 people, what is the probability that at least one person would have type B blood
Answer #1

Probability of blood type B = 0.10

Probability that a person does not have blood type B = 1-0.10 = 0.90

Probability that at least one person would have type B blood = 1- P(X = 0)

= 1 - 0.90*0.90*0.90 = 0.271